# Foundation Models For Seismic Data Processing: An Extensive Review

### v1

@@ Abstract
Seismic processing plays a crucial role in transforming raw data into high-quality subsurface images, pivotal for various geoscience applications. Despite its importance, traditional seismic processing techniques face challenges such as noisy and damaged data and the reliance on manual, time-consuming workflows. The emergence of deep learning approaches has introduced effective and user-friendly alternatives, yet many of these deep learning approaches rely on synthetic datasets and specialized neural networks. Recently, foundation models have gained traction in the seismic domain, due to their success in natural imaging. This paper investigates the application of foundation models in seismic processing on the tasks: demultiple, interpolation, and denoising. It evaluates the impact of different model characteristics, such as pre-training technique and neural network architecture, on performance and efficiency. Rather than proposing a single seismic foundation model, this paper critically examines various natural image foundation models and suggest some promising candidates for future exploration.
@@ 1 Introduction
Seismic processing is essential for converting raw data into high-quality subsurface images. Precise subsurface imaging is vital for various geoscience applications, such as sedimentary and tectonic interpretation, hydrocarbon exploration, reservoir analysis, and geothermal characterization Yilmaz and Doherty (2001). The challenges associated with this process include environmental noise, damaged geophones, and weak low-frequency signals, which degrade data accuracy. Therefore, this transformation involves complex processing steps to enhance the quality of the data. While these techniques improve the quality and accuracy of seismic images, they often rely on time-consuming workflows that approximate the physics behind wave propagation. Additionally, these techniques frequently require manual processes like picking velocities and mute functions, as well as parameter tuning, generally involving iterative trial-and-error adjustments based on the dataset, such as discriminating multiples and primaries in the Radon domain Trad et al. (2003). Therefore, seismic processing requires significant human expertise to extract geological and physical information while mitigating undesired artifacts inherent in the data or introduced by the processing.
@@
In contrast, foundation models are large neural networks that have been pre-trained on vast amounts of data in order to be used in a wide range of downstream tasks. This approach has recently led to numerous breakthroughs across diverse domains, including: ChatGPT Brown et al. (2020) in the natural language processing domain, SAM Kirillov et al. (2023) in image segmentation, Stable Diffusion Rombach et al. (2022) in image generation, SORA Liu et al. (2024) in video generation and more relevantly for seismic processing, image encoders, such as Vision Transformer (ViT) Dosovitskiy et al. (2020), ConvNeXt Liu et al. (2022a) and Swin Transformer (Swin) Liu et al. (2021) in CV.
@@ 1.1 FMs in geophysics
Another image-based approach is the generative seismic foundation model (GSFM) Cheng et al. (2025), which utilizes a diffusion model. In contrast, to the SFMs previously mentioned, which pre-train a feature extractor and then train specific decoders for different downstream tasks, their GSFM is pre-trained and fine-tuned for all their downstream tasks simultaneously. This enables them to share the whole network, not just the image encoder, between downstream tasks.
@@ 2 Methodology
The encoder-decoder architecture, illustrated in Figure 2, serves as the foundation of our framework developed to benchmark a broad variety of FMs.
@@
As defined in Equation (2), the encoder outputs a set of embeddings. Each embedding $E_{s}$ is an element of a three-dimensional space, with $C_{s}$ channels, a height of $H_{s}$ and a width of $W_{s}$. The total number of embeddings $S$ depends on the architecture and number of skip-connections. A neural network without skip-connections only uses one embedding as embedding space and a neural network with skip-connections, such as a UNet, uses multiple embeddings. We define an encoder as non-hierarchical, if all embeddings have the same spatial dimensions, as defined in Equation (4). In contrast, a hierarchical model has embeddings at varying resolutions.
@@
As illustrated in Figure 2, our encoder-decoder architecture uses four embeddings. The choice of four stages is motivated by the fact that most hierarchical models, that we compared in this study, have four distinct stages.
An FM is a model that has been pre-trained on a vast dataset and is then used for several distinct downstream tasks. One possible pre-training method is self-supervised learning as outlined in Equation (5). We chose self-supervised learning as the example, because is the most probable pre-training method for an SFM and was already used by Sheng et al. (2024), Pham et al. (2025) and Sansal et al. (2025). It should be noted though, that other pre-training methods exist. We will discuss some of them later in Section 2.1.1.
@@
For self-supervised learning, as outlined in Equation (5) the mapping changes from $X\mapsto Y$ to $X^{M}\mapsto X$, with $X$ being our data, such as natural images or seismic field data and $X^{M}$ being a randomly masked version of $X$. Consequently, $X^{*}$ is the prediction of the encoder-decoder network ($X^{*}\coloneqq\mathcal{D}_{\theta^{PT}_{D}}(\mathcal{E}_{\theta^{PT}_{E}}(X^{M}))$). Additionally, $L_{ss}$ is a pixel based loss, such as $\ell_{1}$ or $\ell_{2}$, with the specific loss depending on the implementation. It is important to note, that while an encoder-decoder network is used during pre-training only the encoder is the FM, as defined in Equation (6). Therefore, the decoder is kept as small as possible ($|\theta^{PT}_{D}|\ll|\theta^{PT}_{E}$|) shifting more responsibility to the encoder to create more meaningful features.
@@
Pre-training with self-supervision can be done on real data. However, downstream training, as outlined in Equation (7), often relies on supervised learning, which, in seismic processing, involves the use of synthetic data, with $X$ representing the input, $Y$ the label, and $Y^{*}$ the prediction of the model ($Y^{*}\coloneqq\mathcal{D}_{\theta_{D}}(FM_{\theta_{E}|\theta^{PT}_{E}}(X))$). Similar to self-supervised learning, during supervised learning any pixel based loss can be used for $L_{s}$. However, based on some initial experiments and the findings by Zhao et al. (2017) we decided on the $\ell_{1}$ for our downstream training.
@@
The reason for choosing a UNet structure is its effectiveness in pixel-level regression tasks, such as the seismic processing tasks we consider. This robustness was demonstrated multiple times, for example by Fernandez et al. (2025) for demultiple, by Fernandez et al. (2022b) and Chai et al. (2020) for interpolation, by Zhang et al. (2019) for denoising and by Cheng et al. (2025) who used a UNet as basis for their GSFM. While the focus of this study are FMs and therefore the encoder, some additional details and investigations about the decoder and its design can be found in Appendix A.2.
@@ 2.1 Models
Table 1: Overview and categorization of the investigated models. For the pre-training strategy SL stands for supervised learning, SSL for self-supervised learning, SSLD for self-supervised learning with distillation, Semi-S for semi-supervised learning, MTL for multitask learning and CL for contrastive learning. The color next to the model denotes the model archetype, and in conjunction with the shape next to the pre-training strategy, this combination constitutes the legend for the majority of the figures in Section 3.
@@ 2.1.1 Pre-Training Methods
The investigated models vary not only in their architecture but also in their pre-training method. Observed methods include supervised, self-supervised, self-supervised with distillation, semi-supervised, multitask and contrastive learning. Short introductions to each method are provided below, with further details available in their references.
@@
Self-supervised learning, in contrast to supervised-learning, utilizes unlabeled image datasets. In this context, a concrete form of self-supervised learning called Masked Image Modeling (MIM) Bao et al. (2022) is used. During training part of the image is masked, and the network has to recreate the image or predict the missing patches. The advantage is that unlabeled image datasets can be used. Furthermore, recreating the image is a more involved task than classification, which leads to more generalizable features.
@@
Self-supervised learning with distillation, as used by Oquab et al. (2024) uses a teacher network, typically significantly larger than the student network, to teach the student to predict the outputs from the teacher. These outputs are a probability distribution over the target classes. By learning to predict the complete probability distribution of the teacher instead of just the correct label, the knowledge representation of the teacher model is distilled into the student model Hinton et al. (2015). This is important because the smaller model may have insufficient capacity to learn this representation on their own.
@@
Semi-supervised training refers to the Noisy Student Training introduced in Xie et al. (2020). It has three main steps: training a teacher on labeled images, using the teacher to generate pseudo labels on unlabeled data, training a student on a combination of labeled and unlabeled data. This process is iterated by switching the student to the teacher and training a new student. This approach not only delivers better accuracy but also better generalization than supervised training. Additionally, similar to self-supervised training this method can benefit from vast unlabeled datasets. In contrast to self-supervised learning with distillation, only the pseudo labels are predicted and not the whole probability distribution of the teacher. A similar iterative approach, but without training multiple networks, was used by Cheng et al. (2025) to shift their network from synthetic data to real data.
@@
Multitask learning references the training employed in Xiao et al. (2024) and combines image-level, pixel-level and semantic tasks. Image-level tasks involve comprehending the overall context of an image and the semantic relationships, with tasks like image classification, captioning, and visual question answering. Pixel level tasks focus on object and entity localization within images and understanding their relationships and spatial context, including object detection and segmentation. Semantic tasks require a detailed understanding of both text and image to align image regions with corresponding text phrases, challenging the model to grasp local details and semantic contexts. In contrast, to the pre-training methods mentioned before, multitask learning is not image to image based, but requires a combination of image and text prompt as input and delivers a text prompt as output. One disadvantage of this approach is that it requires an annotated dataset.
@@
Finally, contrastive learning Radford et al. (2021) aims to learn a shared feature space for images and text. To achieve this a text encoder and image encoder are trained together to minimize the cosine similarity between correct (text, image) pairings and maximize the distance between incorrect pairings. One disadvantage of this approach, similar to multitask learning, is that it requires an annotated dataset. On the other hand this approach enables zero- and few-shot classification, not limited by the number of classes used during training. Furthermore, models pre-trained with contrastive learning enabled breakthroughs in image retrieval and search. Not only is it possible to search images based on text, but also text based on images.
@@
The majority of the models examined in this study have a corresponding pre-training method, with exceptions being ConvNeXt and Swin. Both are available pre-trained with supervision and self-supervision (both models have a V2 version of the model for self-supervision with only minimal architectural changes). Additionally, ConvNeXt (in its V1 variant) is available with contrastive learning as pre-training strategy. Furthermore while DEIT3, DINO and EVA02 are distinct models, fundamentally they are all a ViT with their biggest difference being their pre-training method. The corresponding pre-training method for each tested model is declared in Table 1.
@@ 2.2 Downstream Training Strategies
The first strategy, dubbed frozen encoder employs FMs pre-trained on natural images directly (freezes the encoder) and only trains the decoder to adapt the output features of the FM for seismic processing tasks, as defined in Equation (8). As described in the equation, freezing the encoder means that the parameters of the encoder stay untouched and are not further optimized. Analogue to Equation (7), $X$ is our input, $Y$ are our labels, $Y^{*}$ are the predictions of the encoder-decoder network and $L_{s}$ is the $\ell_{1}$ loss.
@@
In most of the experiments, the strategy frozen encoder was employed because it is significantly cheaper to compute. This is because the encoder, which corresponds to upwards of $80\%$ of the whole encoder-decoder model, is frozen and not trained. Our hypothesis is that the transfer learning capability of the network is indicative of the performance of the network for seismic pre-training. Furthermore, we aim to investigate the impact of natural image pre-training. Natural image datasets, unlike curated seismic data sets, are extensive. Furthermore, due to solely training the decoder the training takes hours instead of days, thereby enabling the assessment of a greater number of FMs.
@@ 3 Results
In order to draw comparisons between the various FMs, it is necessary to consider a number of different dimensions. The primary and most significant of these is the architecture, which can be roughly divided into hierarchical and non-hierarchical models, as well as transformer-, convolution- and hybrid-based approaches. The categorization of the models was discussed in Section 2.1. Another differentiating factor is the pre-training strategy, which is predominantly interesting from a transfer learning perspective and to a lesser extent to the training of a seismic processing FM. This is due to the fact that the majority of pre-training strategies are not applicable to the seismic processing domain. A more relevant aspect, is the pre-training dataset size used, as this provides insight into the amount of data required to train an FM for seismic processing. The final dimensions that will be examined in this study are model size and inference time, as well as the size of the embedding space. These are crucial because an SFM must be efficient enough to handle the substantial amount of data involved in seismic processing.
@@ 3.1.1 Experimental Setup and Data
For the interpolation and denoising tasks, we use the open source 2007 BP Anisotropic Velocity Benchmark Shah (2007) dataset, consisting of 1641 shot gathers of size $800\times 1151$. Due to the prominence of the first break and the extensive empty sections above it in the 2007 BP Anisotropic Velocity Benchmark, random cuts of the size $224\times 224$ are applied below the first break. This is done, in order to ensure that each gather contains meaningful seismic information.
For all three applications a random sign change with a $50\%$ chance of occurring is applied. Furthermore, for interpolation and denoising $50\%$ of the time the gather is flipped horizontally. While the additive inversion and horizontal flips are applied after the z-scoring and normalizing, random noise is added before. The added random noise is independently sampled for each gather and defined by its maximum strength, rate of occurrence and distribution. For interpolation and demultiple similar noise is added, with a maximum strength of $25\%$ of the maximum absolute value of the gather, a $50\%$ rate of occurrence and a uniform distribution. This noise is added to the input as well as to the labels. In contrast, the noise for denoising is only applied to the input data. Its maximum strength is $50\%$ of the maximum absolute amplitude of the gather, it is always applied and stems from a Gaussian distribution.
Additionally, the gathers are z-scored per gather and then normalized between $-1$ and $1$. This is done to match the value ranges of the seismic data to the natural image ones used during pre-training.
3.1.2 Impact of natural image pre-training
Figure 4: A comparison of the downstream training strategies for the Swin model family. Each plot in the figure depicts the PSNR achieved by the Swin model trained with the three distinct training strategies on a growing downstream task dataset for a specific downstream task. Plot a) shows the results for the demultiple task, plot b) for the interpolation task and plot c) for the denoising task. The observed ranking of these downstream training strategies is fine-tuned encoder $\geq$ non-pre-trained encoder $\geq$ frozen encoder, independently of the downstream task and dataset size.
Figure 4 presents a comparison of the downstream training strategies, outlined in the 2.2 section, for the Swin model family. The Swin model family was chosen, because it is the best performing model across all three tasks, that we tested, as we will show in the following section. Each plot in the figure depicts the PSNR achieved by the Swin model trained with the three distinct training strategies on a growing downstream task dataset. From left to right, the plots show the results for the demultiple, interpolation and denoising task. The observed ranking of these downstream training strategies is fine-tuned encoder $\geq$ non-pre-trained encoder $\geq$ frozen encoder, independently of the downstream task and dataset size. This demonstrates that pre-training on natural images is always comparable to or superior to random initialization. The latter holds particularly for lower data regimes, but even for larger datasets, such as the demultiple dataset. This suggests that during natural image pre-training, the model may develop capabilities that extend beyond those learnable from the specific downstream task.
Furthermore, the figure demonstrates the validity of our assumption, that the performance of the strategy frozen encoder is correlated with that of the other two strategies. Consequently, it can be inferred that better performance using the strategy frozen encoder should be reflected by improved performance using the other two strategies. Therefore, the results in the subsequent sections are from models trained with the strategy frozen encoder, except for our baselines. This is because the strategy frozen encoder, while performing worse than the other two, is significantly cheaper to compute. As baselines, we employed a Swin and a ConvNeXt model (both in their V1 variant) trained with strategy non-pre-trained encoder, which is the dominant strategy used in seismic processing.
3.1.3 Impact of the FM Architecture
Figure 5 presents three graphs, which plot the PSNR for the different seismic processing tasks against each other. While only the PSNR is displayed here, the structural similarity index measure (SSIM) and mean squared error (MSE) were also recorded, but not included in the plots as they are highly correlated with the PSNR, as evidenced in Appendix A.1. The legend encompasses the backbone archetype and the pre-training strategy. The former pertains to a model family, for instance, all ConvNeXt models are under the CONVNEXT archetype, irrespective of the pre-training strategy and version (V1 and V2). Furthermore, all model sizes (TINY, SMALL, BASE) fall under the same archetype, as seen later in Figure 10. The latter were introduced earlier. The only addition is NO_PRE_TRAINING which indicates that the model was trained end-to-end for this specific downstream task without a pre-trained encoder (strategy non-pre-trained encoder). These models serve as a baseline for comparison. All other models were trained using strategy frozen encoder.
Figure 5: Comparison of different FMs on the various seismic processing tasks. Each plot contrasts the PSNR of two downstream tasks. Plot a) shows the results for the interpolation task contrasted against the demultiple task, plot b) for the denoising task against the demultiple task and plot c) for the denoising task against the interpolation task. The horizontal lines delineate the PSNRs of the hierarchical and non-hierarchical models to demonstrate the performance gap between them. Swin models perform the best the across all three tasks, followed by Hiera, CAFormer and ConvNeXt depending on the downstream task.
As indicated above, the results demonstrate that non-hierarchical models perform significantly worse than their hierarchical counterparts. To visually emphasize this discrepancy, horizontal lines have been drawn in Figure 5, delineating the PSNRs of these two model types. The exception to that, are the models ViTamin and EfficientNet on the demultiple task, where they perform significantly worse than the other hierarchical models and similar to the non-hierarchical models, with Dual Attention Vision Transformer (DaViT) even outperforming them. Otherwise, the best performance of a non-hierarchical model is again DaViT this time for denoising, where it is significantly better than the other non-hierarchical models, but still achieved a PSNR of 1.1 $\mathrm{dB}$ less than the worst hierarchical model achieved.
Overall the best performing models are Swin models, followed by Hiera, CAFormer or ConvNeXt models depending on the downstream task. While for demultiple and interpolation there is a significant gap between them and our baselines the gap narrows for denoising, especially when only considering the Swin models. Interestingly, while the Swin archetype always outperforms the ConvNeXt archetype, the ConvNeXt baseline is slightly better in demultiple and denoising than the Swin baseline. This may indicate that the Swin models learn more generalizable image features during pre-training and lend itself better to transfer learning.
3.1.4 Pre-Training Strategy and Downstream Task Correlation
Interestingly, in our findings the optimal pre-training strategies differ between tasks. The interpolation and denoising tasks both perform better with self-supervised pre-training when available, as is the case with Swin and ConvNeXt models. Demultiple, conversely, performs better with supervised or contrastive pre-training for Swin and ConvNeXt respectively. One thing to keep in mind, is that while the dataset for self-supervised and supervised pre-training was the same, the dataset for contrastive pre-training was significantly bigger, as we will show in following section.
Downstream task performance is highly correlated especially for the hierarchical models, which is also similar to the findings of Goldblum et al. (2023). The specific Pearson’s correlations are shown in Appendix A.1. While the non-hierarchical models perform similarly for interpolation, there is a small difference between the results for demultiple and a significant difference between the results for denoising. This correlation between downstream tasks also supports the creation of an SFM, because an SFM should collectively maximize its performance on all downstream tasks.
3.1.5 Combined Downstream Task Performance

### v2

@@ Abstract
Seismic processing plays a crucial role in transforming raw data into high-quality subsurface images, pivotal for various geoscience applications. Despite its importance, traditional seismic processing techniques face challenges such as noisy and damaged data and the reliance on manual, time-consuming workflows. The emergence of deep learning approaches has introduced effective and user-friendly alternatives, yet many of these deep learning approaches rely on synthetic datasets and specialized neural networks. Recently, foundation models have gained traction in the seismic domain, due to their success in the natural image domain. Therefore, we investigate the application of natural image foundation models on the three seismic processing tasks: demultiple, interpolation, and denoising. We evaluate the impact of different model characteristics, such as pre-training technique and neural network architecture, on performance and efficiency. Rather than proposing a single seismic foundation model, we critically examine various natural image foundation models and suggest some promising candidates for future exploration.
@@ 1 Introduction
Seismic processing is essential for converting raw data into high-quality subsurface images. Precise subsurface imaging is vital for various geoscience applications, such as sedimentary and tectonic interpretation, hydrocarbon exploration, reservoir analysis, and geothermal characterization Yilmaz and Doherty (2001). The challenges associated with this process include among others environmental noise, damaged geophones, and weak low-frequency signals, which degrade data accuracy. Therefore, this transformation involves complex processing steps, relying on time-consuming workflows that approximate the physics behind wave propagation, to enhance the quality of the data. Additionally, these techniques frequently require manual processes like picking velocities and mute functions, as well as parameter tuning, generally involving iterative adjustments based on the dataset. Therefore, seismic processing requires significant human expertise to extract geological and physical information while mitigating undesired artifacts inherent in the data or introduced by the processing.
@@
In contrast, foundation models are large neural networks that have been pre-trained on vast amounts of data in order to be used in a wide range of downstream tasks. This approach has recently led to numerous breakthroughs across diverse domains, including: ChatGPT Brown et al. (2020) in the natural language processing domain, SAM Kirillov et al. (2023) in image segmentation, Stable Diffusion Rombach et al. (2022) in image generation, SORA Liu et al. (2024) in video generation and image encoders, such as Vision Transformer (ViT) Dosovitskiy et al. (2020), ConvNeXt Liu et al. (2022a) and Swin Transformer (Swin) Liu et al. (2021).
@@ 1.1 FMs in geophysics
Another image-based approach is the generative seismic foundation model (GSFM) by Cheng et al. (2025), which utilizes a diffusion model. In contrast, to the SFMs previously mentioned, their GSFM is pre-trained and fine-tuned for all their downstream tasks simultaneously. This enables them to share the whole network, not just the image encoder, between downstream tasks.
@@ 2 Methodology
An FM is a neural network pre-trained on a vast dataset that serves several distinct downstream tasks. The encoder-decoder architecture, illustrated in Figure 2, serves as the basis of our framework developed to benchmark a broad variety of FMs.
@@
As defined in Equation (2), the encoder outputs a set of embeddings. Each embedding $E_{s}$ is an element of a three-dimensional space, with $C_{s}$ channels, a height of $H_{s}$ and a width of $W_{s}$. The total number of embeddings $S$ depends on the architecture and number of skip-connections. A neural network without skip-connections only uses one embedding and a neural network with skip-connections, such as a UNet, uses multiple embeddings. We define an encoder as non-hierarchical, if all embeddings have the same spatial dimensions, as defined in Equation (4). In contrast, a hierarchical model has embeddings at varying resolutions.
@@
The choice of four embeddings ($S=4$), as illustrated in Figure 2, is motivated by the fact that most hierarchical models, that we compared in this study, have exactly four distinct stages.
Several pre-training methods exist. One possible pre-training method is self-supervised learning as outlined in Equation (5). We chose self-supervised learning as the example, because it is the most likely pre-training method for an SFM and was already used by Sheng et al. (2024), Pham et al. (2025) and Sansal et al. (2025). In this context, self-supervised learning refers to Masked Image Modeling (MIM) Bao et al. (2022), which is the most common technique within the self-supervised learning paradigm.
@@
For self-supervised learning the mapping changes from $X\mapsto Y$ to $X^{M}\mapsto X$, where $X$ represents the original data, and $X^{M}$ denotes a randomly masked version of $X$. Consequently, $X^{*}$ is the prediction of the encoder-decoder network ($X^{*}\coloneqq\mathcal{D}_{\theta^{PT}_{D}}(\mathcal{E}_{\theta^{PT}_{E}}(X^{M}))$). Additionally, $L_{ss}$ is a pixel based loss, such as $\ell_{1}(X,Y)=||X-Y||_{1}$ or $\ell_{2}(X,Y)=||X-Y||_{2}$, with the specific loss depending on the implementation. It is important to note, that while an encoder-decoder network is used during pre-training, only the encoder is the FM, as defined in Equation (6). Therefore, the decoder is kept as small as possible ($|\theta^{PT}_{D}|\ll|\theta^{PT}_{E}$|) shifting more responsibility to the encoder in order to create more meaningful features.
@@
Pre-training with self-supervision can be done on real data. However, downstream training, as outlined in Equation (7), often relies on supervised learning, with $X$ representing the input, $Y$ the label, and $Y^{*}$ the prediction of the model ($Y^{*}\coloneqq\mathcal{D}_{\theta_{D}}(FM_{\theta_{E}|\theta^{PT}_{E}}(X))$). Similar to self-supervised learning, during supervised learning any pixel based loss can be used for $L_{s}$. However, based on some initial experiments and the findings by Zhao et al. (2017) we decided on the $\ell_{1}$ loss for our downstream training.
@@
The reason for choosing a UNet structure is its effectiveness in pixel-level regression tasks, such as the seismic processing tasks we consider. This robustness has been demonstrated multiple times, for example by Fernandez et al. (2025) for demultiple, by Fernandez et al. (2022b) and Chai et al. (2020) for interpolation, by Zhang et al. (2019) for denoising and by Cheng et al. (2025) who used a UNet as basis for their GSFM, among others.
@@ 2.1 Models
Table 1: Overview and categorization of the investigated models. For the pre-training strategy SL stands for supervised learning, SSL for self-supervised learning, SSLD for self-supervised learning with distillation, Semi-S for semi-supervised learning, MTL for multitask learning and CL for contrastive learning. The color next to the model denotes the model archetype, and in conjunction with the shape next to the pre-training strategy, this combination constitutes the legend for the majority of the plots in this study.
@@ 2.1.1 Pre-Training Methods
The investigated models vary both in their architecture and in their pre-training method. Observed methods include supervised, self-supervised, self-supervised with distillation, semi-supervised, multitask and contrastive learning. Short introductions to each method are provided below, with further details available in their references.
@@
Self-supervised learning, in contrast to supervised-learning, utilizes unlabeled image datasets. In this context, a concrete form of self-supervised learning called MIM Bao et al. (2022) is used. During training part of the image is masked, and the network has to recreate the image or predict the missing patches. The advantage is that unlabeled image datasets can be used. Furthermore, recreating the image is a more involved task than classification, which leads to more generalizable features He et al. (2021).
@@
Self-supervised learning with distillation Oquab et al. (2024) uses a teacher network, typically significantly larger than the student network, to teach the student to predict the outputs from the teacher. These outputs are a probability distribution over the target classes. By learning to predict the complete probability distribution of the teacher instead of just the correct label, the knowledge representation of the teacher model is distilled into the student model Hinton et al. (2015). This is important because the smaller model may have insufficient capacity to learn this representation on their own.
@@
Semi-supervised training refers to the Noisy Student Training introduced in Xie et al. (2020). It has three main steps: training a teacher on labeled images, using the teacher to generate pseudo labels on unlabeled data, training a student on a combination of labeled and unlabeled data. This process is iterated by switching the student to the teacher and training a new student. This approach not only delivers better accuracy but also better generalization than supervised training Xie et al. (2020). Additionally, similar to self-supervised training this method can benefit from vast unlabeled datasets. In contrast to self-supervised learning with distillation, only the pseudo labels are predicted and not the whole probability distribution of the teacher. A similar iterative approach, but without training multiple networks, was used by Cheng et al. (2025) to shift their network from synthetic data to real data.
@@
Multitask learning references the training employed in Xiao et al. (2024) and combines image-level, pixel-level and semantic tasks. Image-level tasks involve comprehending the overall context of an image and the semantic relationships, with tasks like image classification, captioning, and visual question answering. Pixel level tasks focus on object and entity localization within images and understanding their relationships and spatial context, including object detection and segmentation. Semantic tasks require a detailed understanding of both text and image to align image regions with corresponding text phrases, challenging the model to grasp local details and semantic contexts. In contrast to the pre-training methods mentioned before, multitask learning is not image-to-image based, but requires a combination of image and text prompts as input and delivers a text prompt as output. Additionally, it requires an annotated dataset.
@@
Finally, contrastive learning Radford et al. (2021) aims to learn a shared feature space for images and text. To achieve this a text encoder and image encoder are trained together to minimize the cosine similarity between correct (text, image) pairings and maximize the distance between incorrect pairings. However, similar to multitask learning, it requires an annotated dataset. On the other hand, this approach enables zero- and few-shot classification, not limited by the number of classes used during training.
@@
The majority of the models examined in this study have a corresponding pre-training method, with the exceptions of ConvNeXt and Swin. Both are available pre-trained with supervision (V1) and self-supervision (V2), with only minimal architectural changes between them. Additionally, ConvNeXt (in its V1 variant) is available with contrastive learning as pre-training strategy. Furthermore while DEIT3, DINO and EVA02 are distinct models, fundamentally they are all ViTs with their the difference being their pre-training method. The corresponding pre-training method, as well as the architecture, for each tested model is summarized in Table 1.
@@ 2.2 Downstream Training Strategies
The first strategy, dubbed frozen encoder employs FMs pre-trained on natural images directly (freezes the encoder) and only trains the decoder to adapt the output features of the FM for seismic processing tasks. As described in Equation (8), freezing the encoder means that the parameters of the encoder stay untouched and are not further optimized. Analogue to Equation (7), $X$ is our input, $Y$ are our labels, $Y^{*}$ are the predictions of the encoder-decoder network and $L_{s}$ is the $\ell_{1}$ loss.
@@
In most of the experiments, the strategy frozen encoder was employed because it is significantly cheaper to compute. This is because the encoder, which corresponds to upwards of $80\%$ of the whole encoder-decoder model, is frozen and not trained. Our hypothesis is that the transfer learning capability of the network is indicative of the performance of the network for seismic pre-training. Furthermore, we aim to investigate the impact of natural image pre-training. Natural image datasets, unlike curated seismic data sets, are extensive and freely available. Furthermore, due to solely training the decoder the training takes hours instead of days, thereby enabling the assessment of a greater number of FMs.
@@ 3 Results
In order to draw comparisons between the various FMs, it is necessary to consider a number of different dimensions. The primary and most significant of these is the architecture, which can be roughly divided into hierarchical and non-hierarchical models, as well as transformer-, convolution- and hybrid-based approaches. The classification of the models was presented earlier in Table 1. Another differentiating factor is the pre-training strategy, which is predominantly interesting from a transfer learning perspective and to a lesser extent to the training of a seismic processing FM. This is due to the fact that the majority of pre-training strategies are not applicable to the seismic processing domain. A more relevant aspect, is the pre-training dataset size used, as this provides insight into the amount of data required to train an FM for seismic processing. The final dimensions that will be examined in this study are model size and inference time. These are crucial because an SFM must be efficient enough to handle the substantial amount of data involved in seismic processing.
@@ 3.1.1 Experimental Setup and Data
For the interpolation and denoising tasks, we use the open source 2007 BP Anisotropic Velocity Benchmark Shah (2007) dataset, consisting of 1641 shot gathers of size $800\times 1151$. Due to the prominence of the first break and the extensive empty sections above it in the 2007 BP Anisotropic Velocity Benchmark, random cuts of the size $224\times 224$ are applied below the first break. This is done, in order to ensure that each gather contains meaningful seismic information. For interpolation, we randomly mask the input gather and for denoising we add random uniform noise to the input gather. Details of further augmentation and pre-processing steps are available in the source code.
In this study, each experiment recorded the peak signal-to-noise ratio (PSNR), structural similarity index measure (SSIM) Wang et al. (2004), and mean squared error (MSE) metrics. However, due to the high correlation among these metrics and a high correlation across the three downstream tasks only the combined SSIM (($SSIM_{combined}=SSIM_{Demultiple}+SSIM_{Interpolation}+SSIM_{Denoise}$)) is displayed. The SSIM was chosen for the combined score as it has a known upper limit of one, ensuring an equal weighting of the various downstream tasks and providing a theoretical upper bound for performance.
3.1.2 Impact of the FM Architecture and Pre-Training Method
Figure 4 presents the combined SSIM for each model relative to the parameters of the FM utilized as the image encoder. The legend includes the backbone archetype and pre-training methods. The archetype refers to a model family; for example, all ConvNeXt models belong to the CONVNEXT archetype, regardless of pre-training strategy or version (V1 and V2). Besides the pre-training methods outlined previously, the legend also includes NO_PRE_TRAINING, indicating models trained with downstream training strategy non-pre-trained encoder, see Equation(10). These models serve as a baseline for comparison, while all other models were trained using the frozen encoder strategy, see Equation (8). On the left, all tested models are shown; on the right, only models with a combined SSIM above 2.5 are displayed, excluding non-hierarchical models, as well as EfficientNet and ViTamin. The top-performing model is Swin (V2) pre-trained via self-supervision, followed by Swin (V1) pre-trained via supervision. The third-best is ConvNeXt (V2) pre-trained via self-supervision, and the fourth-best is CAFormer, closely followed by Hiera. Interestingly, CAFormer performs well despite being trained via supervision. For both Swin and ConvNeXt, the self-supervised variants surpass their supervised counterparts, suggesting that pre-training CAFormer via self-supervision could significantly enhance results, potentially exceeding ConvNeXt (V2).